OpenStudy (anonymous):

which ordered pair is a solution of the equation y=-5x+4 (2,14) (5,-29) (3,-19) (5,-21)

jigglypuff314 (jigglypuff314):

when given a point (x, y) plug in the value of x for x and the value of y for y in the given equation and solve, and if the statement comes out to equal itself then the ordered pair is on the line so for example given (2, 14) you plug into y=-5x+4 so 14 = -5(2) + 4 14 = -10 + 4 14 = -6 14 does not equal -6 so the answer is not A

OpenStudy (anonymous):

who said it was a ?

jigglypuff314 (jigglypuff314):

I was just proving a point so that you would have an idea of how to get the answer

jigglypuff314 (jigglypuff314):

so given (5, -21) you plug into y = -5x + 4 -21 = -5(5) + 4 -21 = -25 + 4 -21 = -21